OpenStudy (anonymous):

|3y+7| = |2y-1| for all real y

OpenStudy (anonymous):

The book says it is equivalent to either 3y+7=2y-1 or 3y+7=-(2y-1).

OpenStudy (anonymous):

Why isn't it ±(3y+7)=±(2y-1)

OpenStudy (anonymous):

I know the answer, that's not what I care about.

OpenStudy (anonymous):

you mean 3y+7=2y-1 and 3y+7=-(2y-1)?

OpenStudy (anonymous):

not "or"

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

(I'm pretty sure it is or, but I'd like to see where you will go with this.)

OpenStudy (anonymous):

never mind. It's more easy than that

OpenStudy (anonymous):

if the absolute values are equal it means that the actual values are either same sign either oposite sign. That's it

OpenStudy (anonymous):

@DanielHendrycks

OpenStudy (anonymous):

Understood.

OpenStudy (anonymous):

happy to help